5.52 fluid mechanics streer 9ed

for two hydraulic machines to be homologous, these must be
(A) be geometrically similar, (B) have the same discharge coefficient when viewed as a hole, Q1/(A1(2gH1)0.5) = Q2/(A2(2gH2)0.5)and (c) have the same ratio of peripheral velocity to fluid velocity, wD/(Q/A). Show that the scale relationships can be expressed as Q/ND3= constant and H/(ND)2= constant, N is the speed of rotation.
Answer #1

Equating head coefficients

D2 DI

N2 N1 1200 240 5

Hi H2 (ND) (NZD2) H, N D H N D 2 2

H2 (5/4)2 = 1.5625- 300

H2=468.75 meter

Equating the flow coefficients Q ND N D Q2-N2 D2 Q ND or

5 02. 600L/S || 을 || = .078125

Q2=46.875 L/SEC


answered by: ANURANJAN SARSAM